Three collie puppies together weigh 8.6kg. Harris is 0.3kg heavier than Jura. Jura is 100g heavier than Lewis. What does each puppy weigh?
Answer by TimothyLamb(4379)

h + j + L = 8.6
h = j + 0.3
j = L + 0.1
L = j - 0.1
---
h + j + L = 8.6
j + 0.3 + j + j - 0.1 = 8.6
3j + 0.2 = 8.6
3j = 8.4
---
j = 2.8 kg
L = 2.7 kg
h = 3.1 kg
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php